OpenStudy (anonymous):

a jar has three black marbles and three white marbles.If you draw three marbles replacing each marble before drawing the next one,what is the probability that you will draw at least one black marble?

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

well, we can use complemen probability... 1 - P(3Black) = 1 - (3C3/6C3) = 1 - 1/20 = 19/20

OpenStudy (anonymous):

this is not the answer

OpenStudy (raden):

are there the options

OpenStudy (anonymous):

yup 1/2 3/8 5/8 7/8

OpenStudy (raden):

does even takes marbles one by one or 3 marbles direct ?

OpenStudy (anonymous):

i think one by one

OpenStudy (raden):

oh ya, evens take marble was return... it means P(1B,1W,1W) +P(1B,1B,1W) + P(1B,1B,1B) = 3/6*3/6*3/6 + 3/6*3/6*3/6 + 3/6*3/6*3/6 = 1/8 + 1/8 + 1/8 = 3/8